A chemist has 800 grams of a dye solution that is 20% of it original strength. How much dye must be added to increase the strength to 50%?
Found 2 solutions by checkley71, Paul:
Answer by checkley71(8403)

800*.2+x=(800+x).5 or 160+x=400+.5x or .5x=400-160 or .5x=240 or x=240/.5 or
x=480 grams added.
proof 160+480=(800+480).5 or 640=1280*.5 or 640=640

Answer by Paul(988)

800(0.2)+x=0.5(x+800)
160+x=0.5+400
0.5=240
x=480
Hence, about 480 should be added
PAul.